# A multiplexed control architecture for superconducting qubits with row-column addressing

### 06 Mar 2024 v1

@@ Abstract
In state-of-the-art superconducting quantum processors, each qubit is controlled by at least one control line that delivers control pulses generated at room temperature to qubits at millikelvin temperatures. This strategy has been successfully applied to control hundreds of qubits but is unlikely to be scalable to control thousands of qubits, let alone millions or even billions of qubits needed in fault-tolerance quantum computing. The reason for this is due to the wiring challenge, the number of accommodated control lines is limited by factors, such as the cooling power and physical space of the cryogenic system, the control footprint area at the qubit chip level, and so on. Here, we introduce a multiplexed control architecture for superconducting qubits with two types of shared control lines, row and column lines, providing an efficient approach for parallel controlling $N$ qubits with $O(\sqrt{N})$ control lines. With the combination of the two-type shared lines, unique pairs of control pulses are delivered to qubits on each row-column intersection, enabling parallel qubit addressing. Of particular concern here is that, unlike traditional gate schemes, both single- and two-qubit gates are implemented with pairs of control pulses. Considering the inherent parallelism and the control limitations, the integration of the architecture into quantum computing systems should be tailored as much as possible to the specific properties of the quantum circuits to be executed. As such, the architecture could be scalable for executing structured quantum circuits, such as quantum error correction circuits.
@@ I Introduction
To perform quantum computing with noisy qubits for solving valuable problems that are intractable for classical computing, quantum error correction (QEC) is widely recognized as the ultimate solution Terhal2015 ; Dennis2002 . To date, among various candidate quantum systems for realizing QEC, superconducting qubits have been demonstrated as a leading one Campbell2024 ; Krinner2022 ; Zhao2022 ; Sundaresan2023 ; Acharya2023a . In a superconducting quantum processor with tens of qubits, a recent experiment has shown that QEC begins to suppress logical errors with increasing system sizes Acharya2023a . Nevertheless, to fully utilize the power of QEC, such as surface code Dennis2002 ; Fowler2012 , in quantum computing, it is generally believed that high-fidelity control over millions or even billions of qubits is required Fowler2012 ; Gidney2021 .
@@
In state-of-the-art superconducting quantum processors, as shown in Fig. 1(a), each qubit is controlled by at least one control line, which delivers dedicated control pulses generated at room temperature to qubits operating at millikelvin temperatures Chen2018 ; Krinner2019 ; Krantz2019 . Thus, for controlling a $N$-qubit system, both the numbers of control lines (input-output connections, IOs) at room temperature $P_{RT}$ and in the dilution refrigerator $P_{cryo}$ are determined by the IO terminals of the qubit chip $P_{chip}$, which is scale linearly with the qubit number $N$, see Fig. 1(b), leading to $P_{RT}=P_{cryo}=P_{chip}\sim O(N)$ Frankea2019 . This independent control strategy can provide great flexibility in controlling qubits Krantz2019 and is applied to achieve low gate errors across small-scale systems with tens or hundreds of qubits. However, given the heating loads from the lines Krinner2019 , the non-negligible feature sizes of the cables or IO terminals, and at least one digital-to-analog converter (DAC) per line for qubit control Krantz2019 , it is not scalable to large-scale systems. This is the direct manifestation of the wiring challenge in scaling-up quantum processors Vandersypen2017 ; Frankea2019 ; Reilly2019 ; Martinis2020 , i.e., the number of control lines that can be managed is limited by factors, such as the cooling power and geometric space of the cryogenic system, the control footprint area at the qubit chip level, the overhead of classical electronics and so on. To facilitate a more intuitive understanding of such a challenge, there is an insightful comparison, to deliver a superconducting quantum processor with just a few thousand qubits, the number of control lines or IOs is adequate for controlling one billion transistors in state-of-the-art classical processors Vandersypen2017 . In this sense, the wiring challenges need to be addressed before a large-scale superconducting quantum processor becomes feasible.
@@
To alleviate the wiring challenge in superconducting quantum systems, strategies which aim to reduce the control lines running from room temperature to cryogenic temperature, i.e., achieving $P_{RT}<P_{Cryo}$, have been explored Asaad2016 ; Lecocq2021 ; McDermott2018 ; Bardin2019 ; Dijk2020 ; Chakraborty2022 ; Takeuchi2023a ; Takeuchi2023b ; Acharya2023 . One of the most widely studied approaches is the implementation of cryogenic control electronics operated at 4 K McDermott2018 ; Bardin2019 ; Dijk2020 ; Chakraborty2022 ; Takeuchi2023a ; Takeuchi2023b or 10 mK Takeuchi2023a ; Takeuchi2023b ; Acharya2023 , but implementing successful qubit control at scale while achieving ultra-low-power dissipation remains challenging. Additionally, on-chip control electronics have also been explored to reduce the chip IO terminals McDermott2018 ; Liu2023 . However, as the control footprint area at the chip level is limited by qubit size, besides the heating dissipation and the newly added noises, the on-chip integration can increase the complexity of wire routing, especially when scaling up. Therefore, a common view is that only taking these strategies alone is unlikely to address the wiring challenge Vandersypen2017 .
@@
Besides the above top-down approaches, one alternative that has emerged from taking a bottom-up perspective is to exploit multiplexed qubit control with shared lines to reduce the line overhead. Such a perspective is extensively explored for semiconductor spin qubits Vandersypen2017 ; Hill2015 ; Veldhorst2017 ; Li2018 and superconducting quantum annealing processors Johnson2010 ; Bunyk2014 . For superconducting qubits, the most successful demonstration is the frequency-multiplexed qubit readout Jerger2011 ; Jerger2012 ; Chen2012 , where a single feedline is shared by several readout circuitries, allowing simultaneous readout of several qubits. However, exploring the multiplexing for qubit control is still in its infancy Li2021 ; Bejanin2022 ; Zhao2023 ; Shi2023 . Unlike the independent qubit control, the multiplexed control will degrade the control flexibility. Thus, to realize quantum computing, generally, new control overhead or hardware components shall be introduced Zhao2023 ; Shi2023 . Most importantly, further demonstrations of these new features that are compatible with high-fidelity qubit control are required.
@@
In this work, we propose a multiplexed control architecture for superconducting qubits with row-column addressing, which could in principle enable parallel control of $N$ qubits with $O(\sqrt{N})$ control lines. With the combination of two types of shared control lines, i.e., row and column lines, unique pairs of control pulses are delivered to qubits or couplers on each row-column intersection, allowing parallel single- or two-qubit gate operations in qubit lattices. Accordingly, we present various single- and two-qubit schemes that are both compatible with the row-column addressing and the existing superconducting qubit technologies (with no new hardware components). Moreover, we show that while the control flexibility is compromised, the inherent parallelism allows the architecture to be well-suitable for executing structured quantum circuits, which consist of layers of parallel single- or two-qubit gates across qubit lattices, such as QEC circuits.
@@
The paper is organized as follows. In Sec. II, we first provide an overview of the multiplexed control architecture. Then, in QEC with surface code, we illustrate that the integration of the architecture into quantum computing systems should be tailored to the specific quantum circuits. In Sec. III, we give detailed descriptions of the single- and two-qubit gate schemes that support row-column addressing. In Sec. IV, we discuss the challenges to be faced when scaling up. Finally, in Sec. V, we provide a summary of our study.
Figure 1: Schematic (not to scale) of a typical superconducting quantum computing system. (a) Qubits at millikelvin temperatures are controlled by delivering control pulses generated by classical electronics at room temperature through the cables. The number of control lines (input-output connections, IOs) at room temperature $P_{RT}$ and the number of cables in the dilution refrigerator $P_{cryo}$ are determined by the number of input-output (IO) terminals of the chip $P_{chip}$, leading to $P_{RT}=P_{cryo}=P_{chip}$. (b) With the independent control strategy, $P_{chip}\sim O(N)$, where $N$ is the qubit number. This strategy allows maximum flexibility in controlling qubits, i.e., each qubit (see the qubit highlighted in light orange) can be individually addressed without affecting any other qubits. (c) Within the multiplexed control architecture, by applying appropriate pulses to each of these control lines simultaneously, unique pairs of control pulses are delivered to qubits that are located at the intersection of row lines and column lines, allowing parallel qubit addressing. Hence, the IO terminals $P_{chip}$ scales with $\sqrt{N}$. Here, each qubit (see the qubit highlighted in light orange) or any subgroup of qubits cannot be individually addressed without affecting all other idle qubits.
@@ II The multiplexed control architecture with row-column addressing
Figure 2: Multiplexed qubit control for QEC with surface code. (a) A typical stabilizer measurement circuit in surface code using CZ gates, in which the single-qubit gate is further decomposed into two $\sqrt{\textmd{X}}$ gates and three Z gates. During the QEC cycle, interleaved single- and two-qubit gates are implemented and at the end of the cycle, syndrome qubits are measured first and then reset while data qubits are idle. (b) Two sets of shared control lines are employed for independent addressing of data qubits and syndrome qubits. This is adopted to ensure the compatibility between the multiplexed control architecture and the QEC circuits including dynamical decoupling sequences for suppressing qubit idling errors and two-qubit gates with one-qubit driving. Here the shared control lines can deliver microwave pulses or flux pulses for controlling qubits. (c) In QEC with surface code, the order of two-qubit gates depends on the type of syndrome qubits, Z-type or X-type, leading to four different patterns for the sequence of two-qubit gates in the QEC cycle. Hence, for QEC based on two-qubit gates with driven couplers, besides shared lines for qubit addressing, an additional four sets of shared lines are required for selectively choosing gate patterns.
@@
For controlling two-dimensional (2D) square qubit lattices, Figure 1(c) schematically shows the multiplexed control architecture, which comprises two types of shared control lines, row lines and column lines. The qubits are located on the intersections of the two-type lines, thus each qubit is connected (coupled) to a unique pair of control lines, enabling spatial addressing of qubits. When applying appropriate pulses to each of these shared control lines simultaneously, each qubit can be driven by a unique pair of control pulses through the associated row and column lines, e.g., see Fig. 1(c). As we will discuss in the next section, when the pair of control pulses satisfies certain conditions, single- or two-qubit gate operations can be realized. Thus, this control strategy can be well suited for performing parallel addressing, allowing simultaneous single- or two-qubit gates across the qubit lattice. Meanwhile, as control lines are shared by qubits, here, each qubit or any subgroup of qubits cannot be selectively addressed without affecting all other idle (inactive) qubits. For example, as shown in Fig. 1(c), when one considers addressing the qubit highlighted in light orange, two pulses are delivered to the associated row and column lines. Thus, idle qubits that are controlled by the same row or column control lines, can be driven by one of the two pulses, causing idling qubit errors.
@@
Evidently, compared to the independent control, there exist two features in controlling qubits. The first feature is that while the multiplexed control can degrade the control flexibility, it is efficient in examining the wiring challenge, allowing parallel control of $N$ qubits with $O(\sqrt{N})$ lines. The second one is that, unlike traditional gate schemes, both single- and two-qubit gates shall be implemented with a pair of control pulses, thus supporting the row-column addressing. Correspondingly, this leads to two issues to be resolved in the architecture, the degradation in control flexibility can affect the usefulness of the multiplexed control in quantum computing, and gate operations supporting the row-column addressing need to be verified. In the following, we aim to first tackle the former one. We will start with a discussion of being agnostic to the physical details of qubits and gate schemes, which will be discussed in the next section.
@@
Although the control flexibility is compromised, the inherent parallelism of the multiplexed control could allow for executing structured quantum circuits, which consist of layers of parallel single- and two-qubit gates across qubit systems (note here that by using the isomorphous waveform technique introduced in Ref. Han2024 , an arbitrary quantum circuit composed of modular single- and two-qubit gates can be compiled into the structured quantum circuits). However, to be practical and useful, the multiplexed control should be tailored to be compatible with the specific properties of the structured circuits.
@@
Here, we consider the realization of a QEC circuit with the multiplexed control scheme. Figure 2(a) shows the typical quantum circuit for stabilizer measurements in surface code using CZ gates Acharya2023a , in which the single-qubit gate is decomposed into two $\sqrt{\textmd{X}}$ gates and three Z gates McKay2017 (the reason for adopting such decomposition will be discussed in the next section). During the QEC cycle, interleaved single- and two-qubit gates are implemented and at the end of the cycle, syndrome qubits are measured first and then reset while data qubits are idle. To make the control scheme compatible with the QEC circuit, Figure 2(b) shows the multiplexed control architecture comprising two sets of shared row-column lines for independently addressing date qubits and syndrome qubits. This tailored setting is adopted for two reasons: (i) during the syndrome measurement and reset, dynamical decoupling that comprises a sequence of single-qubit gates is generally applied to data qubits for suppressing idling errors (the idling is currently the dominant error) Acharya2023a ; (ii) in certain two-qubit gate schemes, only one of the two qubits, e.g., the syndrome qubit shown in the bottom inset of Fig. 2(c), is driven by control pulses Chow2011 ; Caldwell2018 .
Additionally, in certain two-qubit gate schemes, control pulses are only applied to couplers McKay2016 ; Mundada2019 ; Han2020 ; Kubo2023 , e.g., see the upper inset of Fig. 2(c). Meanwhile, in stabilizer measurement circuits of surface code, the order of two-qubit gates depends on the measured syndromes, Z- or X. This gives rise to four different patterns for the sequence of two-qubit gates in QEC cycles Terhal2015 ; Dennis2002 , as shown in Fig. 2(c). Therefore, here, besides the shared lines for qubit addressing, an additional four sets of shared lines are required for selectively choosing the gate pattern and addressing couplers in the subgroup of the whole lattice.
Before leaving this section, note that while here we only focus on QEC circuits, we expect that the multiplexed control can also be tailored for logical operations. This is supported by the following three observations. First, logical operations, such as using lattice surgery Horsman2012 ; Fowler2018 , in essence, involve different patterns of stabilizer measurements Fowler2018 , which themselves are compatible with the multiplexed control. Second, a more tailored setting can be introduced for patterns of stabilizers that are incompatible with the above setting and the added line overhead can scale with $\sqrt{N}$. Third, while selectively addressing a few subgroups of qubits, i.e., selecting patterns of stabilizers for logical operations, can result in qubit idling errors, the errors, e.g., phase errors due to off-resonance drives, could be compensated in subsequent QEC cycles.
@@
Figure 3: Schematic illustration of single-qubit gate schemes with two control pulses. (a) Single-qubit gates applied to frequency-tunable transmon qubits. During single-qubit gate operations, a flux pulse and a microwave pulse are simultaneously delivered to the qubit through the Z line and the XY line, respectively. The static DC flux biases are needed for biasing the qubits at their idle points and are realized by using an on-chip programmable magnetic memory, e.g., a $\Phi$-DAC. (ST1) illustrates the gate scheme based on tuning the qubit on-resonance with the microwave drive pulse according to a baseband flux pulse. Individual tuning of the coupling between the Z line and the qubit is achieved by a tunable rf-SQUID coupler, which is biased by a $\Phi$-DAC. (ST2) shows the gate scheme that utilizes the combination of the parametric drive and the microwave drive. Due to the nonlinear dependence of the qubit frequency on the flux bias, the parametric-driven qubit can have a series of sideband frequency components. When one of such sidebands is on-resonance with the microwave drive, qubit control can be achieved. (b) Single-qubit gates for fixed-frequency transmon qubits. During single-qubit gates, a pair of microwave drive pulses are simultaneously delivered to the qubit through the two XY lines. (SF1) depicts the gate schemes using the three-wave mixing process mediated by a flux qubit. When the sum-frequency or difference-frequency of the two microwave drives is on-resonance with the qubit, the qubit can be driven by an effective on-resonance drive. (SF2) displays the gate scheme that uses one off-resonance microwave drive to control a qubit dressed by a second off-resonance drive.

### 20 Mar 2024 v2

@@ Abstract
In state-of-the-art superconducting quantum processors, each qubit is controlled by at least one control line that delivers control pulses generated at room temperature to qubits operating at millikelvin temperatures. While this strategy has been successfully applied to control hundreds of qubits, it is unlikely to be scalable to control thousands of qubits, let alone millions or even billions of qubits needed in fault-tolerance quantum computing. The primary obstacle lies in the wiring challenge, wherein the number of accommodated control lines is limited by factors, such as the cooling power, physical space of the cryogenic system, the control footprint area at the qubit chip level, and so on. Here, we introduce a multiplexed control architecture for superconducting qubits with two types of shared control lines, row and column lines, providing an efficient approach for parallel controlling $N$ qubits with $O(\sqrt{N})$ control lines. With the combination of the two-type shared lines, unique pairs of control pulses are delivered to qubits at each row-column intersection, enabling parallel qubit addressing. Of particular concern here is that, unlike traditional gate schemes, both single- and two-qubit gates are implemented with pairs of control pulses. Considering the inherent parallelism and the control limitations, the integration of the architecture into quantum computing systems should be tailored as much as possible to the specific properties of the quantum circuits to be executed. As such, the architecture could be scalable for executing structured quantum circuits, such as quantum error correction circuits.
@@ I Introduction
To perform quantum computing with noisy qubits for solving valuable problems that are intractable for classical computing, quantum error correction (QEC) is widely recognized as the ultimate solution Terhal2015 ; Dennis2002 . To date, among the various candidate quantum systems for realizing QEC, superconducting qubits have been demonstrated as a leading one Campbell2024 ; Krinner2022 ; Zhao2022 ; Sundaresan2023 ; Acharya2023a . In a superconducting quantum processor with tens of qubits, a recent experiment has shown that QEC begins to suppress logical errors with increasing system sizes Acharya2023a . Nevertheless, to fully utilize the power of QEC, such as surface code Dennis2002 ; Fowler2012 , in quantum computing, it is generally believed that high-fidelity control over millions or even billions of qubits is required Fowler2012 ; Gidney2021 .
@@
In state-of-the-art superconducting quantum processors, shown in Fig. 1(a), each qubit is controlled by at least one control line, which delivers dedicated control pulses generated at room temperature to qubits operating at millikelvin temperatures Chen2018 ; Krinner2019 ; Krantz2019 . For controlling a $N$-qubit system, the number of control lines (input-output connections, IOs) at room temperature ($P_{RT}$) and in the dilution refrigerator ($P_{cryo}$) is dictated by the IO terminals of the qubit chip ($P_{chip}$), scaling linearly with the qubit number $N$, see Fig. 1(b), resulting in $P_{RT}=P_{cryo}=P_{chip}\sim O(N)$ Frankea2019 . This independent control strategy offers great flexibility in controlling qubits Krantz2019 and is applied to achieve low gate errors across small-scale systems comprising tens or hundreds of qubits. However, the scalability of this approach to larger systems is hindered by challenges such as heating loads from the control lines Krinner2019 , non-negligible feature sizes of cables or IO terminals, and the requirement of at least one digital-to-analog converter (DAC) per line for qubit control Krantz2019 . These limitations highlight the wiring challenge associated with scaling up quantum processors Vandersypen2017 ; Frankea2019 ; Reilly2019 ; Martinis2020 , where the number of manageable control lines is limited by factors including cooling power, geometric constraints of the cryogenic system, control footprint area at the qubit chip level, and the overhead of classical electronics. To facilitate a more intuitive understanding of this challenge, consider that to control a superconducting quantum processor with a few thousand qubits, the number of required control lines or IOs would be comparable to that needed to manage one billion transistors in state-of-the-art classical processors Vandersypen2017 . In this sense, the wiring challenges need to be addressed before a large-scale superconducting quantum processor becomes feasible.
@@
To alleviate the wiring challenge in superconducting quantum systems, various strategies which aim to reduce the control lines running from room temperature to cryogenic temperature, i.e., achieving $P_{RT}<P_{Cryo}$, have been explored Asaad2016 ; Lecocq2021 ; McDermott2018 ; Bardin2019 ; Dijk2020 ; Chakraborty2022 ; Takeuchi2023a ; Takeuchi2023b ; Acharya2023 . One of the most widely studied approaches is the implementation of cryogenic control electronics operated at 4 K McDermott2018 ; Bardin2019 ; Dijk2020 ; Chakraborty2022 ; Takeuchi2023a ; Takeuchi2023b or 10 mK Takeuchi2023a ; Takeuchi2023b ; Acharya2023 , but implementing successful qubit control at scale while achieving ultra-low-power dissipation remains challenging. Additionally, on-chip control electronics have also been explored to reduce the chip IO terminals McDermott2018 ; Liu2023 . However, as the control footprint area at the chip level is limited by qubit size, besides the heating dissipation and the newly added noises, the on-chip integration can increase the complexity of wire routing, especially when scaling up. Therefore, a common view is that only taking these strategies alone is unlikely to address the wiring challenge Vandersypen2017 .
@@
Besides the above top-down approaches, one alternative that has emerged from taking a bottom-up perspective involves leveraging multiplexed qubit control with shared lines to reduce line overhead. This perspective has been extensively explored for semiconductor spin qubits Vandersypen2017 ; Hill2015 ; Veldhorst2017 ; Li2018 and superconducting quantum annealing processors Johnson2010 ; Bunyk2014 . For superconducting qubits, the most successful demonstration is the frequency-multiplexed qubit readout Jerger2011 ; Jerger2012 ; Chen2012 , where a single feedline is shared by several readout circuitries, enabling simultaneous readout of several qubits. However, exploring the multiplexing for qubit control is still in its infancy Li2021 ; Bejanin2022 ; Zhao2023 ; Shi2023 . Unlike the independent qubit control, the multiplexed control will degrade the control flexibility. Thus, to realize quantum computing, generally, new control overhead or hardware components shall be introduced Zhao2023 ; Shi2023 . Most importantly, further demonstrations of these new features that are compatible with high-fidelity qubit control are required.
@@
In this work, we propose a multiplexed control architecture for superconducting qubits with row-column addressing, offering the potential for parallel control of $N$ qubits with $O(\sqrt{N})$ control lines. By incorporating both row and column shared control lines, unique pairs of control pulses are delivered to qubits or couplers at each row-column intersection, allowing parallel single- or two-qubit gate operations in qubit lattices. Accordingly, we present various single- and two-qubit schemes that are both compatible with the row-column addressing and the existing superconducting qubit technologies (with no new hardware components). We further show that while the control flexibility is compromised, the inherent parallelism of the architecture makes it well-suited for executing structured quantum circuits, which comprise layers of parallel single- or two-qubit gates across qubit lattices, such as QEC circuits.
@@
The paper is organized as follows. In Sec. II, we first provide an overview of the multiplexed control architecture. Then, in QEC with surface code, we illustrate that the integration of the architecture into quantum computing systems should be tailored to specific quantum circuits. In Sec. III, we give detailed descriptions of the single- and two-qubit gate schemes that support row-column addressing. In Sec. IV, we discuss the challenges to be faced when scaling up. Finally, in Sec. V, we provide a summary of our study.
Figure 1: Schematic (not to scale) of a typical superconducting quantum computing system. (a) Qubits at millikelvin temperatures are controlled using control pulses generated by classical electronics at room temperature and delivered through cables. The number of control lines (input-output connections, IOs) at room temperature $P_{RT}$ and the number of cables in the dilution refrigerator $P_{cryo}$ are determined by the number of input-output (IO) terminals of the chip $P_{chip}$, leading to $P_{RT}=P_{cryo}=P_{chip}$. (b) With the independent control, $P_{chip}\sim O(N)$, where $N$ is the qubit number. This strategy offers maximum flexibility in qubit control, enabling individual addressing without affecting other qubits (see the qubit highlighted in light orange). (c) In the multiplexed control architecture, applying appropriate pulses to control lines simultaneously allows unique pairs of control pulses to be delivered to qubits at the intersection of row lines and column lines, enabling parallel qubit addressing. Hence, the IO terminals $P_{chip}$ scales with $\sqrt{N}$. Here, each qubit (see the qubit highlighted in light orange) or any subgroup of qubits cannot be individually addressed without affecting all other idle qubits.
@@ II The multiplexed control architecture with row-column addressing
Figure 2: Multiplexed qubit control for QEC with surface code. (a) A typical stabilizer measurement circuit in surface code using CZ gates, with the single-qubit gate decomposed into two $\sqrt{\textmd{X}}$ gates and three Z gates. During the QEC cycle, interleaved single- and two-qubit gates are implemented and at the end of the cycle, syndrome qubits are measured first and then reset while data qubits are idle. (b) Two sets of shared control lines are employed for independent addressing of data qubits and syndrome qubits. This setup ensures compatibility between the multiplexed control architecture and QEC circuits, which involve dynamical decoupling sequences to reduce qubit idling errors and two-qubit gates with one-qubit driving. The shared control lines can deliver microwave or flux pulses for qubit control. (c) In QEC with surface code, the sequence of two-qubit gates depends on the type of syndrome qubits (Z-type or X-type), resulting in four different patterns for the sequence of two-qubit gates in the QEC cycle. Hence, for QEC based on two-qubit gates with driven couplers, besides shared lines for qubit addressing, an extra four sets of shared lines are required for selectively choosing gate patterns.
@@
For controlling two-dimensional (2D) square qubit lattices, Figure 1(c) schematically shows the multiplexed control architecture, which comprises two types of shared control lines, row lines and column lines. The qubits are located on the intersections of the two-type lines, with qubit connected (coupled) to a distinct pair of control lines, enabling spatial qubit addressing. By simultaneously applying appropriate pulses to the shared control lines, each qubit can be driven by a unique pair of control pulses through the associated row and column lines, as depicted in Fig. 1(c). As we will discuss in the next section, when the pair of control pulses satisfies certain conditions, single- or two-qubit gate operations can be realized. Thus, this control strategy can be well suited for performing parallel addressing, allowing simultaneous single- or two-qubit gates across the qubit lattice.
However, due to the shared nature of the control lines among qubits, selective addressing of individual qubits or specific subgroups (i.e., for implementing non-structured quantum circuits) without impacting all other idle (inactive) qubits is challenging Tan2024 . For example, as shown in Fig. 1(c), when one considers addressing the qubit highlighted in light orange, two pulses are delivered to the associated row and column lines. Consequently, idle qubits coupled to the same row or column control lines can be driven by one of the two shared pulses, causing idling qubit errors. Similarly, one can envision a scenario in which fine-tuned gate pulses are simultaneously applied to all the shared lines. This allows gate operations on select qubits to be executed successfully. However, the non-selective (inactive) qubits are also driven by pairs of pulses, potentially leading to idle gate errors. Note that as each qubit is controlled by a unique pair of gate pulses, this allows for tailored design of pulses to implement distinct gates on individual qubits. For instance, target gates can be applied to the active qubits while identity operations can be performed on the idle qubits. However, this approach generally results in the introduction of additional overheads in gate calibration.
@@
In comparison to the independent control, there exist two features in controlling qubits. The first feature involves the trade-off between multiplexed control and control flexibility. While multiplexed control can reduce control flexibility, it effectively addresses wiring challenges by enabling parallel control of $N$ qubits with $O(\sqrt{N})$ lines. The second feature pertains to the requirement for both single- and two-qubit gates to be executed using a pair of control pulses, thereby supporting row-column addressing. Consequently, two key issues emerge within this architecture: the impact of reduced control flexibility on the utility of multiplexed control in quantum computing, and the validation of gate operations that support row-column addressing. In the subsequent discussion, we aim to first tackle the former one. We will start with a discussion of being agnostic to the physical details of qubits and gate schemes, which will be elaborated in the next section.
@@
Despite the compromise in control flexibility, the inherent parallelism of the multiplexed control could allow for executing structured quantum circuits, which consist of layers of parallel single- and two-qubit gates across qubit systems (note that by using the isomorphous waveform technique introduced in Ref. Han2024 , an arbitrary quantum circuit comprising modular single- and two-qubit gates can be compiled into structured quantum circuits). Nevertheless, for practicality and utility, the multiplexed control should be tailored to be compatible with the specific properties of the structured circuits.
@@
Here, we consider the realization of a QEC circuit with the multiplexed control scheme. Figure 2(a) shows the typical quantum circuit for stabilizer measurements in surface code using CZ gates Acharya2023a , where the single-qubit gate (including the identity gate) is decomposed into two $\sqrt{\textmd{X}}$ gates and three Z gates McKay2017 (the rationale behind this decomposition will be addressed in the subsequent section). Throughout the QEC cycle, interleaved single- and two-qubit gates are implemented and at the end of the cycle, syndrome qubits are measured first and then reset while data qubits are idle. To ensure compatibility with the QEC circuit, Figure 2(b) presents the multiplexed control architecture comprising two sets of shared row-column lines for independently addressing date qubits and syndrome qubits. This tailored setting is adopted for two reasons: (i) during the syndrome measurement and reset, dynamical decoupling comprising a sequence of single-qubit gates is generally applied to data qubits for suppressing idling errors (currently the predominant error source) Acharya2023a ; (ii) in certain two-qubit gate schemes, only one of the qubits, such as the syndrome qubit shown in Fig. 2(c), is driven by control pulses Chow2011 ; Caldwell2018 .
@@
Figure 3: Schematic illustration of single-qubit gate schemes with two control pulses. (a) Single-qubit gates applied to frequency-tunable transmon qubits. During gate operations, a flux pulse and a microwave pulse are simultaneously delivered to the qubit through the Z line and the XY line, respectively. Static dc flux biases are needed for biasing the qubits at their idle points and are realized by using an on-chip programmable magnetic memory such as $\Phi$-DAC. (ST1) illustrates the gate scheme based on tuning the qubit on-resonance with the microwave drive pulse using a baseband flux pulse. Individual tuning of the coupling between the Z line and the qubit is enabled by a tunable rf-SQUID coupler biased by a $\Phi$-DAC. (ST2) shows the gate scheme that utilizes the combination of parametric drive and microwave drive. Due to the nonlinear dependence of the qubit frequency on the flux bias, the parametric-driven qubit can have a series of sideband frequency components. When one of such sidebands is on-resonance with the microwave drive, qubit control can be achieved. (b) Single-qubit gates for fixed-frequency transmon qubits. During gates, a pair of microwave drive pulses are simultaneously delivered to the qubit through the two XY lines. (SF1) depicts the gate scheme using the three-wave mixing process mediated by a flux qubit. When the sum-frequency or difference-frequency of the two drives is on-resonance with the qubit, the qubit can be controlled by an effective on-resonance drive. (SF2) displays the gate scheme that uses one off-resonance microwave drive to control a qubit dressed by a second off-resonance drive.
Additionally, as depicted in Fig. 2(c), specific two-qubit gate schemes involve applying control pulses solely to couplers McKay2016 ; Mundada2019 ; Han2020 ; Kubo2023 . Meanwhile, in stabilizer measurement circuits of the surface code, the order of two-qubit gates depends on the syndromes being measured, either Z or X. This gives rise to four different patterns for the sequence of two-qubit gates in QEC cycles Terhal2015 ; Dennis2002 , as shown in Fig. 2(c). Therefore, to minimize any aforementioned idle errors and calibration overhead, besides the shared lines for qubit addressing, an extra four sets of shared lines can be introduced for selectively choosing the gate pattern and addressing couplers.
Before leaving this section, note that although we only focus on QEC circuits, there is potential for adapting the multiplexed control scheme for logical operations as well. This is supported by the following three key observations. First, logical operations, such as those based on lattice surgery Horsman2012 , in essence, involve different patterns of stabilizer measurements that are compatible with the multiplexed control. Second, a more tailored setting can be introduced for patterns of stabilizers that are incompatible with the above setting and the added line overhead can scale with $\sqrt{N}$. Third, while selectively addressing specific qubit subgroups to accommodate stabilizer patterns for logical operations may cause qubit idling errors, such as phase errors due to off-resonance drives, this could be compensated in subsequent QEC cycles, albeit potentially leading to an increase in circuit depth.